# Demystifying Invariant Effectiveness for Securing Smart Contracts

### v1

@@ 5.2. Invocation Tree Analysis
Invocation tree analysis extracts invariant-related data exclusively from the invocation tree, sufficient to collect data for access control, time lock, gas control, re-entrancy, oracle, and money flow invariants. For instance, for the sender opcode, its results can be extracted from the invocation tree by searching for parent node of the node of target contract. Moreover, the invocation tree can also be used to identify locations of re-entrancy by capturing the nested and recursive calls to the target contract. Oracle values are also obtained by traversing the invocation tree to locate calls to the oracle. For money flow invariants, the required values can also be directly read from function calls of transfering Ether/ERC20.
@@ 5.3. Type Inference
Type inference in Trace2Inv infers the type of storage slot when it is accessed. It is essential for extracting data relevant to both special storage and data flow invariants, as accurately decoding storage accesses with the correct type is necessary for these invariant categories. Decoding storage slots is straightforward when they are listed in the contract’s storage layout, which is typically the case for special storage invariants. However, the challenge arises with complex data structures like mapping, often used in data flow invariants. These structures may reference storage slots not explicitly present in the contract’s storage layout, which instead are computed through operations such as sha3 and arithmetic functions. To solve this issue, Trace2Inv maintain a preimage dictionary to track the key’s computation. When Trace2Inv encounters a sha3 opcode, the mapping positions and slots are recorded. In Solidity, the first $32$-bytes represent the mapping slot, and the next $32$-bytes serve as the mapping position. In Vyper, these roles are reversed. Anytime a 64-byte sha3 hash is encountered, both its hash value and origin are recorded for future reference. When an sload operation has a key that is not present in the storage layout, the preimage dictionary is consulted. We recursively trace back the computation steps until we identify a mapping data structure that is in the storage layout. Utilizing the types of the mapping data structure, we can infer the type of storage slot accessed.
@@ 5.4. EVM-level Dynamic Taint Analysis
Sources External Address Variables balance, extcodesize, extcodecopy, extcodehash,
Execution Context Variables origin, caller, address, codesize, selfbalance, pc, msize, gas
Call Data Variables callvalue, calldataload, calldatasize, calldatacopy
Return Data Variables returndatasize, returndatacopy
Block Variables blockhash, coinbase, timestamp, number, prevrandao, gasprice, gaslimit, chainid,
@@
Bit Level Taint Propagation. Our taint analyzer maintains three distinct taint trackers: the stack tracker, the memory tracker, and the storage tracker. Initially, all these trackers are empty. Each tracker records taint information at the bit level, enabling granular analysis. For instance, sload and sstore opcodes can only read and write $32$-byte chunks, but the taint status is stored for each individual bit. Our propagation of taints follows a set of rules based on prior work (Ghaleb et al., 2022):
@@ 5.6. Trace2Inv Implementation
Trace2Inv’s input, transaction trace data, can be obtained from Ethereum via its API debug_traceTransaction. Within the Trace Parser module, Trace2Inv queries EtherScan (Eth, 2020) to gather additional transaction metadata not directly available in the trace data, such as block number and transaction origin. This metadata provides the execution context crucial for generating certain invariants, like EOA and LU. Additionally, Trace2Inv also queries EtherScan (Eth, 2020) to fetch the contract’s source code. The code is then compiled locally using the appropriate versions of the Solidity (sol, 2018) or Vyper (vyp, 2018) compilers to obtain the contract’s function ABI. With this ABI, Trace2Inv leverages Slither (sli, 2021) to decode function names, their arguments, and return values present in the trace. Within Invariant Related Data Extraction Module, Trace2Inv compiles source code of target smart contract, and reads the compiler’s output to obtain its storage layout. However, some old versions of Solidity and Vyper compilers do not support this functionality. In such cases, Trace2Inv fetches the storage layout from EVM Storage (EVM, 2024).
Several optimizations are incorporated into Trace2Inv implementation to enhance its performance. Firstly, when fetching trace data from an Ethereum archive node, we optimize the process by using batching RPC requests. This significantly reduces the overhead associated with individual API calls. Secondly, we employ parallelization techniques to speed up the process of fetching and parsing trace data. Specifically, multi-processing is used to concurrently handle different segments of trace data, converting them into summaries in a more time-efficient manner. Thirdly, caching is utilized to further optimize performance; the system caches query results from EtherScan, archive nodes, and compilers. This minimizes redundant executions and API calls, thereby accelerating the overall analysis process.
@@ 6. Evaluation
Table 4. Benchmarks. (Exploit: the first exploit transaction during the incident. FL: whether the exploit transaction uses flash loan. Contracts: the victim contracts involved in the exploit. Type: the type of the contract, either interface (I) or payload (P). PL: the programming language of the contract, either Solidity (S) or Vyper (V). History: the number of transactions in the history of the contract up to the exploit transaction.)
@@
    Bridges RoninNetwork keys compromised 22/03/29 624M (Etherscan, 2023u) RoninNetwork I+P S 95345
HarmonyBridge keys compromised 22/06/24 100M (Etherscan, 2023h) HarmonyBridge P S 32149
@@
22/08/01 152M (Etherscan, 2023m) Nomad I+P S 15630
PolyNetwork hash collision 21/08/10 611M (Etherscan, 2023p) PolyNetwork I+P S 44509
    Lending bZx2 oracle manipulation 20/02/18 630K (Etherscan, 2023b) bZx2 I+P S 711
@@
(Etherscan, 2023c) CheeseBank_3 I+P S 557
InverseFi oracle manipulation 22/06/16 1.26M (Etherscan, 2023l) InverseFi I+P S 7590
@@
(Etherscan, 2023e) CreamFi2_4 I+P S 98
RariCapital1 read-only re-entrancy 21/05/09 10M (Etherscan, 2023r) RariCapital1 I+P S 667
@@
(Etherscan, 2023s) RariCapital2_4 I+P S 776
XCarnival logic error 22/06/26 3.87M (Etherscan, 2023z) XCarnival I+P S 342
    Yield-Earning Harvest1 oracle manipulation 20/10/26 33.8M (Etherscan, 2023i) Harvest1 I+P S 2050
Harvest2 oracle manipulation 20/10/26 (Etherscan, 2023j) Harvest2 I+P S 2161
ValueDeFi oracle manipulation 20/11/14 7.4M (Etherscan, 2023w) ValueDeFi I+P S 295
@@
(Etherscan, 2024) Yearn_I I S 26688
VisorFi re-entrancy 21/12/21 8.2M (Etherscan, 2023x) VisorFi I+P S 1693
UmbrellaNetwork underflow bug 22/03/20 700K (Etherscan, 2023v) UmbrellaNetwork I+P S 59
PickleFi access control 20/11/21 20M (Etherscan, 2023o) PickleFi I+P S 5439
    Others Eminence logic error 20/09/29 7M (Etherscan, 2023g) Eminence I+P S 20589
Opyn logic error 20/08/04 371k (Etherscan, 2023n) Opyn I+P S 67
IndexFi logic error 21/10/15 16M (Etherscan, 2023k) IndexFi I+P S 20641
@@
(Etherscan, 2023t) RevestFi_I I S 1463
DODO access control 21/03/08 700K (Etherscan, 2023f) DODO I+P S 42
@@
(Etherscan, 2023a) BeanstalkFarms_I I S 306
Table 4 presents the benchmark dataset in our study. It comprises 27 hacks which cumulatively resulted in financial losses exceeding 2 billion USD. The column FL denotes whether the exploit involves flash loans, which require atomicity for the hack transactions. The column Type denotes the type of victim contracts. For each exploit, two types of victim contracts are manually identified: payload (P) refers to the protocol contract that eventually transfers abnormal amounts of tokens out of the protocol, and interface (I) refers to the contract that is directly invoked by users to initiate this abnormal transfer. The History column signifies the length of the transaction history up to the exploit transaction for each victim contract.
@@ 6.1. RQ1: Effectiveness of Smart Contract Invariants
Results. Table 5 provides a comprehensive summary of the effectiveness evaluation for the invariants applied across various contract categories. The table lists several key metrics: the number of contracts to which each type of invariant is applied, the number of contracts successfully protected by these invariants, the number of hacks blocked, and the Average False Positive rate. Among these, the row # Hacks Blocked stands out as the most crucial metric as it directly measures the capability of each invariant to block financial exploits. The Average FP rate is also an important metric as it quantifies the potential impact on regular users, reflecting the trade-off between security and usability.
@@
For false positives, EOA, OB and TOU have an average FP rate below 0.4%, while GS and DFU also have a low FP rate below 2.4%. This low rate indicates that these invariants have a small impact on regular user transactions, thereby making them practical for real-world deployment. The elevated false positive rates observed for OR, TSU, and TBU are primarily because of the fluctuating nature of oracle values, total supply, and total borrow. Using upper-bound or range-based invariants for these categories could inadvertently block all transactions once these values exceed a certain threshold.
Answer to RQ1: EOA in Access Control, OB in Time Lock, GS in Gas Control, TOU in Money Flow, and DFU in Data Flow shows the highest efficacy in blocking hacks while having a low false positive rate.

### v2

@@ 5.2. Invocation Tree Analysis
Invocation tree analysis extracts invariant-related data from the invocation tree, that is sufficient to collect data for access control, time lock, gas control, re-entrancy, oracle, and money flow invariants. For instance, for the sender opcode, its results is extracted from the invocation tree by searching for parent node of the target contract node. Moreover, the invocation tree is also used to identify locations of re-entrancy by capturing nested and recursive calls to the target contract. Oracle values are also obtained by traversing the invocation tree to locate calls to the oracle. For money flow invariants, the required values are also read from function calls of transferring Ether/ERC20.
@@ 5.3. Type Inference
Type inference in Trace2Inv infers the type of storage slot when it is accessed. It is essential for extracting data relevant to both special storage and data flow invariants, as accurately decoding storage accesses with the correct type is necessary for those invariants. Decoding storage slots is straightforward when they are listed in the contract’s storage layout, which is typically the case for special storage invariants. However, the challenge arises with complex data structures like mapping, often used in data flow invariants. These structures may reference storage slots not explicitly present in the contract’s storage layout, which are instead computed through operations such as sha3 and arithmetic functions. To solve this issue, Trace2Inv maintains a preimage dictionary to track the key’s computation. When we encounter a sha3 opcode, the mapping positions and slots are recorded. In Solidity, the first $32$-bytes represent the mapping slot, and the next $32$-bytes serve as the mapping position. In Vyper, these roles are reversed. Anytime a 64-byte sha3 hash is encountered, both its hash value and origin are recorded. When a sload operation has a key that is not present in the storage layout, the preimage dictionary is consulted. We recursively trace back the computation steps until we identify a mapping data structure that is in the storage layout. Using the types of the mapping data structure, we infer the type of storage slot accessed.
@@ 5.4. EVM-level Dynamic Taint Analysis
Sources External Address balance, extcodesize, extcodecopy, extcodehash
Execution Context origin, caller, address, codesize, selfbalance, pc, msize, gas
Call Data callvalue, calldataload, calldatasize, calldatacopy
Return Data returndatasize, returndatacopy
Block blockhash, coinbase, timestamp, number, prevrandao, gasprice, gaslimit, chainid
@@
Bit Level Taint Propagation. Our taint analyzer maintains three distinct taint trackers: the stack, the memory, and the storage trackers. Initially, all these trackers are empty. Each tracker records taint information at the bit level, enabling granular analysis. For instance, sload and sstore opcodes can only read and write $32$-byte chunks, but the taint status is stored for each individual bit. Our propagation of taints follows a set of rules based on prior work (Ghaleb et al., 2022):
@@ 5.6. Trace2Inv Implementation
Trace2Inv’s input, transaction trace data, can be obtained from Ethereum via its API debug_trace-Transaction. Within the Trace Parser module, Trace2Inv queries EtherScan (Eth, 2020) to gather additional transaction metadata not directly available in the trace data, such as block number and transaction origin. This metadata provides the execution context crucial for generating certain invariants, like EOA and LU. Additionally, Trace2Inv also queries EtherScan (Eth, 2020) to fetch the contract’s source code. The code is then compiled locally using the appropriate versions of the Solidity (sol, 2018) or Vyper (vyp, 2018) compilers to obtain the contract’s function ABI. With this ABI, Trace2Inv leverages Slither (sli, 2021) to decode function names, their arguments, and return values present in the trace. Within Invariant Related Data Extraction Module, Trace2Inv compiles source code of target smart contract, and reads the compiler’s output to obtain its storage layout. However, some old versions of Solidity and Vyper compilers do not support this functionality. In such cases, Trace2Inv fetches the storage layout from EVM Storage (EVM, 2024).
Several optimizations are incorporated into Trace2Inv to enhance its performance. First, when fetching trace data from an Ethereum archive node, we optimize the process by using batching RPC requests. This significantly reduces the overhead associated with individual API calls. Second, we use parallelization techniques to speed up the fetching and parsing trace data. Specifically, multiprocessing is used to concurrently handle different segments of trace data, converting them into summaries in a time-efficient manner. Third, caching is utilized to further optimize performance; the system caches query results from EtherScan, archive nodes, and compilers. This minimizes redundant executions and API calls, thereby accelerating the overall analysis process.
@@ 6. Evaluation
Table 4. Benchmarks+. (Exploit: the first exploit transaction during the incident. FL: whether the exploit transaction uses flash loan. Contracts: the victim contracts involved in the exploit. Type: the type of the contract, either interface (I) or payload (P). PL: the programming language of the contract, either Solidity (S) or Vyper (V). History: the number of transactions in the history of the contract up to the exploit transaction.)
@@
    Bridges RoninNetwork keys compromised 22/03/29 624M (Ron, 2023) RoninNetwork I+P S 95345
HarmonyBridge keys compromised 22/06/24 100M (Har, 2023a) HarmonyBridge P S 32149
@@
22/08/01 152M (Nom, 2023) Nomad I+P S 15630
PolyNetwork hash collision 21/08/10 611M (Pol, 2023) PolyNetwork I+P S 44509
    Lending bZx2 oracle manipulation 20/02/18 630K (bZx, 2023) bZx2 I+P S 711
@@
(Che, 2023) CheeseBank_3 I+P S 557
InverseFi oracle manipulation 22/06/16 1.26M (Inv, 2023) InverseFi I+P S 7590
@@
(Cre, 2023b) CreamFi2_4 I+P S 98
RariCapital1 read-only re-entrancy 21/05/09 10M (Rar, 2023a) RariCapital1 I+P S 667
@@
(Rar, 2023b) RariCapital2_4 I+P S 776
XCarnival logic error 22/06/26 3.87M (XCa, 2023) XCarnival I+P S 342
    Yield-Earning Harvest1 oracle manipulation 20/10/26 33.8M (Har, 2023b) Harvest1 I+P S 2050
Harvest2 oracle manipulation 20/10/26 (Har, 2023c) Harvest2 I+P S 2161
ValueDeFi oracle manipulation 20/11/14 7.4M (Val, 2023) ValueDeFi I+P S 295
@@
(Yea, 2024) Yearn_I I S 26688
VisorFi re-entrancy 21/12/21 8.2M (Vis, 2023) VisorFi I+P S 1693
UmbrellaNetwork underflow bug 22/03/20 700K (Umb, 2023) UmbrellaNetwork I+P S 59
PickleFi access control 20/11/21 20M (Pic, 2023) PickleFi I+P S 5439
    Others Eminence logic error 20/09/29 7M (Emi, 2023) Eminence I+P S 20589
Opyn logic error 20/08/04 371k (Opy, 2023) Opyn I+P S 67
IndexFi logic error 21/10/15 16M (Ind, 2023) IndexFi I+P S 20641
@@
(Rev, 2023) RevestFi_I I S 1463
DODO access control 21/03/08 700K (DOD, 2023) DODO I+P S 42
@@
(Bea, 2023) BeanstalkFarms_I I S 306
Table 4 presents the benchmark dataset in our study. It comprises 27 hacks which resulted in financial losses over $2$ billion USD. The column FL denotes whether the exploit involves flash loans, which require atomicity for the hack transactions. The column Type denotes the type of victim contracts. For each hack, two types of victim contracts are manually identified: payload (P) refers to the contract that eventually transfers abnormal amounts of tokens out of the protocol, and interface (I) refers to the contract that is directly invoked by users to initiate this transfer. The History column gives the length of the transaction history up to the hack transaction for each victim contract.
@@ 6.1. RQ1: Effectiveness of Smart Contract Invariants
Table 5. Complete Results of Invariant Effectiveness Evaluation. (Gray cells indicate the exploit transaction is blocked by corresponding invariant guard(True Positives). The number in the cell represents the number of non-exploit transactions blocked by corresponding invariant guard(False Positives). Symbol - represents that the invariant is by nature not applicable to the benchmark contract. Symbol represents that the transactions in the training set violate the invariant. Symbol $\emptyset$ represents the transactions in the training set do not have enough data points to learn the invariant.)
Access Control Time Lock Gas Control Re Oracle Storage Money Flow Data Flow     Kind Benchmarks     EOA     SO     SM     OO     OM     SB     OB     LU     GS     GC     RE     OR     OD     TSU     TBU     TIU     TOU     TIRU     TORU     MU     CVU     DFU     DFL RoninNetwork 0 0 0 0 - - - - 0 0 0.4 0.1 - 0 0 0 HarmonyBridge 0 0 0 0 0 2.8 2.8 0 - - - - 0 0 0 0 - 0 0 0 Nomad 0 0 0 0 0 0 0 - - - - 0 0 0 0 - - 0 0     Bridges PolyNetwork 0 0 0 0 0 0 0 0 - - - - 0 0 0 0 - 0 - - bZx2 0 0 2.3 0 2.3 0 0 3.3 1.4 1.9 0 $\emptyset$ $\emptyset$ 0.5 0 0 0 - - 0 - 0 0.5 Warp 0 0 0 2.3 0 0 0 6.8 6.8 0 - - - 93.2 0 4.5 0 0 - - 4.5 0 Warp_I 0 0 11.1 0 33.3 - - 0 22.2 33.3 - 88.9 44.4 - - - - - - - - - - CheeseBank_1 0 0 0 0 0 0.5 0.5 4.3 0 7.1 1.6 0 33.2 2.2 1.1 2.2 1.6 - - 3.3 0.5 CheeseBank_2 0 0 0 0 0 0.6 1.1 1.1 0 10.2 2.8 0 7.9 0 0 0 0 - - 0 0 CheeseBank_3 0 0 0 0 0 0 1.8 1.8 0 10.8 1.8 0 7.8 3 0 3 0 - - 3 3.6 InverseFi 0 0 0 0.1 0.1 0 0 0.3 0.1 0 0 0 0 1.9 16.3 1.5 0 1.3 0.1 - - 2 0.2 CreamFi1_1 0 0 $\emptyset$ 0 $\emptyset$ 0 0 $\emptyset$ $\emptyset$ $\emptyset$ 0 - - - - - - - - - - $\emptyset$ $\emptyset$ CreamFi1_2 0.5 0.5 0.1 0.5 0 0.5 0.1 0.3 0 - - 0 0 0 0 0.1 0.2 - 0 0 0 CreamFi2_1 0 0.3 0 0.3 0.3 0 0 0.8 1.1 1.1 0 0 0 2.9 9.5 1.8 0 0 0.8 2.6 - 0.8 0.5 CreamFi2_2 0 0 1.5 0 1.9 0 0 3.7 3.3 2.6 0 0 0 0 0 0 0 0 0.4 2.6 - 0 1.1 CreamFi2_3 3.8 0 9 0 12.8 0 0 7.7 3.8 2.6 0 - - 10.3 0 3.8 0 3.8 0 0 - 3.8 1.3 CreamFi2_4 0 0 0 0 0 0 0 0 0 0 0 - 20.7 0 3.4 0 3.4 3.4 - 3.4 0 RariCapital1 0 0 6 0 4.5 0 0 6.5 1.5 0 0 - - - - 0.5 0 - - - 0.5 0.5 0 RariCapital2_1 0 0 0 0 0 0 0 1.1 1.1 0.5 0 - - 10.9 6 0.5 0.5 0 0 - 0.5 0.5 0 RariCapital2_2 1.8 0 0.4 0 0.4 0 0 1.8 0.4 0.4 0 - - 6.7 2.2 0.9 0 0.9 0.4 1.3 - 0.9 2.7 RariCapital2_3 0 0 0 0 0 0 0 0 0.8 1.7 0 - - 0 0 0 0 0.8 0 1.7 - 0 0 RariCapital2_4 0 0 0 0 0 0 0 0.4 0.4 2.6 0 - - 0 0 0 0 0 0 0.9 - 0 0     Lending XCarnival 0 6.9 5.9 0 0 2 11.8 8.8 0 - - 0 4.9 3.9 1 4.9 28.4 3.9 3.9 1 1 Harvest1 0 0.5 0 0 0 0 0 0.2 0 0 0 29.3 0 1.1 - 0 0.2 0 0.2 - - 2.6 0.2 Harvest2 0 0 0 0 0 0 0 0 0 0 0 34.7 0 0 - 0 0 0.2 0.2 - - 1.2 0.2 ValueDeFi 0 0 0 0 0 3.4 1.1 2.3 0 64.8 0 73.9 - 0 0 0 0 - - 8 1.1 Yearn1 0 0 0 0 0 0 0 0 0 0 - - - - - - - - - - $\emptyset$ $\emptyset$ Yearn1_I 0 0 0 0 0.1 0.1 0 - - - - 0 0.1 0.1 1.8 0 - 4.4 0 VisorFi 0 0 0 0 0 0 0 0 0 - - - - 0 1.2 0 0 - - 4.2 0 UmbrellaNetwork 0 11.8 11.8 0 0 5.9 0 0 0 - - - - 5.9 0 5.9 0 - - 5.9 17.6     Yield-Earning PickleFi 0 0 0 0 0 0 0 0 0.1 0.1 0 - - - - 0 0.1 - - - - 0.1 0.1 Eminence 0 0 0 0 0 0.1 0 0 - - 20.5 - 0 0.5 0 0.1 - - 0.4 0 Opyn 0 0 0 0 0 0 0 0 0 0 0 - - - - - - - - - - $\emptyset$ $\emptyset$ IndexFi 0 0 0 0 0 0 0 0 0.7 0.1 0 - - 6.6 - 0 0 0.1 0 0.4 - 0 0 RevestFi 0 0 0 3.7 0 0 4.1 0.4 1.8 0 - - - - - - - - - - $\emptyset$ $\emptyset$ RevestFi_I 1.4 0 0 0 4.1 0 0 4.6 0.5 2.1 0 - - - - - - - - - - - - DODO 0 0 8.3 0 0 0 0 33.3 25 0 - - 0 - - - - - - - $\emptyset$ $\emptyset$ Punk_1 0 0 $\emptyset$ 0 $\emptyset$ - - 25 0 0 0 - - - - 0 0 - - - - 0 0 Punk_2 0 0 $\emptyset$ 0 0 - - 0 0 0 - - - - 0 0 - - - - 0 0 Punk_3 0 0 $\emptyset$ 0 $\emptyset$ - - 18.2 0 0 0 - - - - 0 0 - - - - 0 0 BeanstalkFarms 0 0 0.2 0 0.4 0 0.2 0.7 0.6 0 - - 31.8 - 0 0.1 0 0 - - 0.8 0.3     Others BeanstalkFarms_I 0 2.2 0 2.2 0 - - 0 0 0 - - - - - - - - - - - - - # Contracts Applied 42 39 22 39 25 33 33 37 41 41 40 11 11 21 16 34 34 28 28 11 7 33 33 # Contracts Protected 23 6 7 5 9 11 13 12 30 23 2 7 6 8 12 9 22 5 22 1 2 23 1 # Hacks Blocked 15 4 6 3 8 9 11 10 18 15 2 5 4 7 6 8 17 5 15 1 2 18 1 Average FP(%) 0.2 0.6 1.8 0.5 2.6 0 0 2.4 2.4 2.6 0 22.3 4.6 7.9 12.6 0.7 0.4 0.8 1.4 1.5 0.7 1.6 0.9
@@
Results. Table 5 provides a detailed breakdown of the effectiveness evaluation for each invariant applied across the $42$ contracts. The table lists the number of contracts on which an invariant is applied, the number of contracts successfully protected by the invariants, the number of hacks blocked, and the false positive rate for each invariant. The row # Hacks Blocked stands out as the most crucial metric as it directly measures the capability of each invariant to block exploits. The average false positive rate is also an important metric as it quantifies the potential impact on regular users, reflecting the trade-off between security and usability.
Table 6 provides a comprehensive summary of the effectiveness evaluation for the invariants applied across various contracts. The table lists several key metrics: the number of contracts on which an invariant is applied, the number of contracts successfully protected by the invariants, the number of hacks blocked, and the average false positive (FP) rate. Among these, the row # Hacks Blocked stands out as the most crucial metric as it directly measures the capability of each invariant to block exploits. The average FP rate is also an important metric as it quantifies the potential impact on regular users, reflecting the trade-off between security and usability.
@@
For false positives, EOA, OB and TOU have an average FP rate below 0.4%, while GS and DFU have a low FP rate below 2.4%. This low rate indicates that these invariants have a small impact on regular user transactions, thereby making them practical for real-world deployment. The elevated false positive rates observed for OR, TSU, and TBU are primarily because of the fluctuating nature of oracle values, total supply, and total borrow. Using upper-bound or range-based invariants for these categories could inadvertently block all transactions once these values exceed a certain threshold.
Answer to RQ1: EOA in Access Control, OB in Time Lock, GS in Gas Control, TOU in Money Flow, and DFU in Data Flow are the most effective in blocking hacks with low false positive rates.